Accumulate ₱ 25,000 for 2 years and 5 months at 6 ½ 3 compounded every two months


1
Expert's answer
2021-04-15T07:40:58-0400

"A=P(1+\frac{r}{n})^{nt}"

"=25000(1+\frac{0.065}{6})^{\frac{29}{12}}"

"=2500\times1.026381757"

"=25659.54392"